11. Which of the following represents a correct chemical formula? Name it.
(a) CaCl
(b) BiPO4
(c) NaSO4
(d) NaS

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) CaCl = Wrong (valency of Ca = 2, Cl = 1) (b) BiPO4 = Correct (valency of Bi = 3, PO4 = 3) (c) NaSO4 = Wrong (valency of Na = 1, SO4 = 2) (d) NaS = Wrong (valency of Na = 1, Sulphide = 2)